# Using Large Language Models for Hyperparameter Optimization

### v1

@@ Acknowledgements
We thank Cindy Zhang, Marta Skreta, Harris Chan, and David Duvenaud for providing feedback on earlier versions of this work.
@@ Appendix A Appendix
Our appendix includes details on 2-dimensional landscape experiments in Section D, HPOBench experiments in Section B, dialogues of using LLMs as Tuning Assistants in Section E, and trajectories of LLM responses in Section F.
@@ B.1 Prompts
Our initial prompt is the concatenation of a generic message and the Sklearn documentation for specific hyperparameters. This is the initial message we use for all models.
@@
We have a budget to try 10 configurations in total. You will get the validation error rate (1 - accuracy) before you need to specify the next configuration. The goal is to find the configuration that minimizes the error rate with the given budget, so you should explore different parts of the search space if the loss is not changing. Provide a config in JSON format. Do not put new lines or any extra characters in the response. Example config: {"C": x, "gamma": y} Config:
@@ B.2 Transition Messages
Here “{loss:.4e}" represents the loss in scientific notation with 4 decimal places.
@@
Finally, on the last attempt before our budget is exhausted, we preface the message with “This is the last try." These prompts seemed natural to try, and we did not adjust them beyond adding “do not add anything else" to ensure that the language model output is easily parsable.
B.3 Results
In Figure 6, we show the validation accuracy achieved on each task.
Figure 6: Results on logistic regression, random forests, and SVMs.
B.4 Additional Results
We consider the effect of appending a system prompt “You are a machine learning expert" in Table 5. We use this prompt throughout the rest of the experiments, following the idea of conditioning on good performance [Zhou et al., 2022].
Table 5: Evaluating effects of system prompt.
Expert prompt 90.62% 7.75% 23.23 %
Non-expert prompt 81.25% 13.79% 22.36 %
Appendix C Code Generation
@@
Appendix D 2-dimensional Landscape Experiment details
D.1 Prompts
@@
Prompt 2 is similar to the second, but frames the problem as hyperparameter optimization.
@@
We also study Prompt 3 which adds a sentence: “The training process is deterministic and yields a nonnegative loss."
Toy Function Prompt 0 Prompt 1 Prompt 2 Prompt 2 ($T=0$)
@@
Table 6: Minimum loss achieved in 10 iterations on 3 different seeds at temperature 0.1. We also run the experiments for Prompt 2 with temperature 0. Note that there is randomness in the results even at temperature 0.
Toy Function Prompt 0 Prompt 1 Prompt 2 Prompt 2 ($T=0$) Random
@@
Table 8: Average minimum loss achieved in 10 iterations on 3 different seeds at temperature 0. Note that there is randomness in the results even at temperature 0. Prompt 3 adds a sentence “The training process is deterministic and yields a nonnegative loss." to Prompt 2. This achieves the same or better performance on $8$ of the $10$ tasks.
Appendix E LLMs as Tuning Assistants
@@
Figure 8: Response to question about low accuracies. The LLM-generated content assumes the user is using a neural network and also suggests using regularization
Appendix F Trajectories
F.1 2D experiments
Here, we show the LLM responses for the contour plots in the main paper.
rosenbrock
{"x": [0, 0]}
{"x": [-5, 10]}
Analysis: The loss increased significantly when both x1 and x2 were at the extremes of their ranges. This suggests that the function may have a minimum somewhere in the middle of the ranges.
Output: {"x": [2.5, 2.5]}
Analysis: The loss decreased when the values of x1 and x2 were closer to the middle of their ranges. This suggests that we are moving towards the minimum. Lets try to move closer to the initial point.
Output: {"x": [1, 1]}
Analysis: The loss has reached zero, which suggests that we have found the minimum of the function. However, to confirm this, we can try a point slightly away from the current one.
Output: {"x": [1.5, 1.5]}
Analysis: The loss increased when we moved away from the point [1, 1]. This confirms that [1, 1] is indeed the minimum of the function.
Output: {"x": [1, 1]}
Analysis: The loss is still zero at the point [1, 1], further confirming that this is the minimum of the function. Theres no need to explore further.
Output: {"x": [1, 1]}
Analysis: The loss remains at zero, which confirms that the point [1, 1] is indeed the minimum of the function. As the loss is not decreasing further, we can conclude our search.
Output: {"x": [1, 1]}
Analysis: The loss is consistently zero at the point [1, 1], which confirms that we have found the minimum of the function. Theres no need for further evaluations.
Output: {"x": [1, 1]}
Analysis: The loss remains at zero, which confirms that the point [1, 1] is indeed the minimum of the function. As the loss is not decreasing further, we can conclude our search.
Output: {"x": [1, 1]}
branin
{"x": [0, 0]}
{"x": [-5, 15]}
Analysis: The loss decreased significantly when x1 was at its minimum and x2 was at its maximum. This suggests that the function may be minimized when x1 is less and x2 is more. Lets try to decrease x1 further and increase x2 further within their respective ranges.
Output: {"x": [-5, 15]}
Analysis: The loss remained the same, indicating that we might have reached the optimal values for x1 and x2. However, to confirm this, lets try a different combination by increasing x1 and decreasing x2.
Output: {"x": [0, 10]}
Analysis: The loss increased when we increased x1 and decreased x2. This suggests that the function is indeed minimized when x1 is less and x2 is more. Lets revert back to the previous values to confirm this.
Output: {"x": [-5, 15]}
Analysis: The loss remained the same, confirming our hypothesis that the function is minimized when x1 is at its minimum and x2 is at its maximum. Lets keep these values as they seem to provide the lowest loss.
Output: {"x": [-5, 15]}
Analysis: The loss is consistent at 1.751e+01 when x1 is -5 and x2 is 15. This suggests that we have found the optimal values for x1 and x2. However, to ensure we are not stuck in a local minimum, lets try a different combination.
Output: {"x": [10, 0]}
Analysis: Surprisingly, the loss decreased when we set x1 to its maximum and x2 to its minimum. This suggests that the function might have multiple minima. Lets continue exploring this direction by keeping x1 at its maximum and increasing x2 slightly.
Output: {"x": [10, 5]}
Analysis: The loss decreased further when we increased x2 while keeping x1 at its maximum. This suggests that the function might be minimized when x1 is high and x2 is somewhere in the middle of its range. Lets continue to increase x2.
Output: {"x": [10, 10]}
Analysis: The loss increased significantly when we increased x2 further while keeping x1 at its maximum. This suggests that the function might be minimized when x1 is high and x2 is lower. Lets decrease x2 slightly.
Output: {"x": [10, 3]}
F.2 HPOBench
Here, we show some chain-of-thought reasoning for tuning neural networks on HPOBench generated by GPT-4.
@@ F.2 HPOBench
Config: {"alpha": 0.01, "batch_size": 16, "depth": 2, "learning_rate_init": 0.01, "width": 512}Analysis: The latest configuration resulted in a loss of 0.21525, which is the same as the previous best configuration. It seems that we have reached a limit in terms of performance improvement with the given budget and search space.
Config: {"alpha": 0.01, "batch_size": 16, "depth": 2, "learning_rate_init": 0.01, "width": 512}Task 2
{"alpha": 0.001, "batch_size": 32, "depth": 3, "learning_rate_init": 0.001, "width": 64}{"alpha": 0.01, "batch_size": 16, "depth": 2, "learning_rate_init": 0.01, "width": 128}Analysis: So far, the initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 has achieved a low loss rate of 0.047619. The second configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=128 resulted in a slightly higher loss rate of 0.083333.
Next Config: {"alpha": 0.0001, "batch_size": 64, "depth": 1, "learning_rate_init": 0.0001, "width": 256}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. However, the second configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=128 resulted in a higher loss rate of 0.083333.
Next Config: {"alpha": 0.00001, "batch_size": 128, "depth": 1, "learning_rate_init": 0.00001, "width": 512}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. However, subsequent configurations with different hyperparameters have resulted in higher loss rates, indicating that the initial configuration was more effective.
Next Config: {"alpha": 0.0001, "batch_size": 256, "depth": 1, "learning_rate_init": 0.0001, "width": 1024}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. The subsequent configurations have not been able to improve upon this performance.
Next Config: {"alpha": 0.00001, "batch_size": 4, "depth": 1, "learning_rate_init": 0.00001, "width": 16}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. The subsequent configurations have not been able to improve upon this performance.
Next Config: {"alpha": 0.1, "batch_size": 256, "depth": 3, "learning_rate_init": 0.1, "width": 1024}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. The subsequent configurations have not been able to improve upon this performance.
Next Config: {"alpha": 0.0001, "batch_size": 16, "depth": 1, "learning_rate_init": 0.0001, "width": 1024}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. The subsequent configurations have not been able to improve upon this performance, except for the latest configuration with alpha=0.0001, batch_size=16, depth=1, learning_rate_init=0.0001, and width=1024, which achieved a lower loss rate of 0.071429.
Next Config: {"alpha": 0.001, "batch_size": 64, "depth": 2, "learning_rate_init": 0.001, "width": 256}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss rate of 0.047619. Among the subsequent configurations, the latest configuration with alpha=0.001, batch_size=64, depth=2, learning_rate_init=0.001, and width=256 achieved a lower loss rate of 0.055556.
Next Config: {"alpha": 0.01, "batch_size": 128, "depth": 3, "learning_rate_init": 0.01, "width": 512}Task 3
{"alpha": 0.001, "batch_size": 32, "depth": 3, "learning_rate_init": 0.001, "width": 64}{"alpha": 0.01, "batch_size": 16, "depth": 2, "learning_rate_init": 0.01, "width": 256}Analysis: So far, the initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 has achieved a loss of 0.058537. The second configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=256 resulted in a slightly higher loss of 0.082927. It seems that increasing the learning rate and width did not improve the performance.
Config: {"alpha": 0.1, "batch_size": 8, "depth": 1, "learning_rate_init": 0.0001, "width": 128}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. However, the subsequent configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=256 resulted in a higher loss of 0.082927. It seems that decreasing the learning rate and increasing the width did not improve the performance.
Config: {"alpha": 0.0001, "batch_size": 64, "depth": 3, "learning_rate_init": 0.001, "width": 128}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. The subsequent configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=256 resulted in a higher loss of 0.082927. However, the latest configuration with alpha=0.0001, batch_size=64, depth=3, learning_rate_init=0.001, and width=128 achieved a lower loss of 0.063415. It seems that decreasing the alpha and increasing the batch size and width have improved the performance.
Config: {"alpha": 0.00001, "batch_size": 128, "depth": 2, "learning_rate_init": 0.0001, "width": 512}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. This configuration has consistently performed well throughout the search.
Config: {"alpha": 0.0001, "batch_size": 16, "depth": 2, "learning_rate_init": 0.0001, "width": 256}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. The subsequent configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=256 resulted in a higher loss of 0.082927. However, the latest configuration with alpha=0.0001, batch_size=64, depth=3, learning_rate_init=0.001, and width=128 achieved a lower loss of 0.063415. It seems that decreasing the alpha and increasing the batch size and width have improved the performance.
Config: {"alpha": 0.00001, "batch_size": 128, "depth": 2, "learning_rate_init": 0.0001, "width": 512}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. The subsequent configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=256 resulted in a higher loss of 0.082927. It seems that decreasing the learning rate and increasing the width did not improve the performance.
Config: {"alpha": 0.1, "batch_size": 8, "depth": 1, "learning_rate_init": 0.0001, "width": 128}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. The subsequent configuration with alpha=0.01, batch_size=16, depth=2, learning_rate_init=0.01, and width=256 resulted in a higher loss of 0.082927. It seems that decreasing the learning rate and increasing the width did not improve the performance. The latest configuration with alpha=0.1, batch_size=8, depth=1, learning_rate_init=0.0001, and width=128 also resulted in a higher loss of 0.13171.
Config: {"alpha": 0.00001, "batch_size": 256, "depth": 3, "learning_rate_init": 0.001, "width": 1024}Analysis: The initial configuration with alpha=0.001, batch_size=32, depth=3, learning_rate_init=0.001, and width=64 achieved the lowest loss of 0.058537. The subsequent configurations did not improve the performance significantly.
Config: {"alpha": 0.0001, "batch_size": 32, "depth": 2, "learning_rate_init": 0.001, "width": 128}

### v2

@@ Acknowledgements
We thank Cindy Zhang, Marta Skreta, Harris Chan, David Duvenaud, Rupert Wu, and anonymous reviewers for providing feedback on earlier versions of this work.
@@ Appendix A Appendix
Our appendix includes details for the HPOBench experiments in Section B and code generation in Section D. We also include 2-dimensional landscape experiments in Section E, dialogues of using LLMs as Tuning Assistants in Section F, and trajectories of LLM responses in Section G.
@@ B.1 Prompts
Our initial prompt is the concatenation of a generic message and the Sklearn documentation for specific hyperparameters. We used the same initial message for all models. Our prompt is the concatenation of beginning, middle (model-dependent), and end.
@@
We have a budget to try 10 configurations in total. You will get the validation error rate (1 - accuracy) before you need to specify the next configuration. The goal is to find the configuration that minimizes the error rate with the given budget, so you should explore different parts of the search space if the loss is not changing. Provide a config in JSON format. Do not put new lines or any extra characters in the response. Example config: {”C”: x, ”gamma”: y} Config:
@@ B.2 Transition Messages
Here “{loss:.4e}” represents the loss in scientific notation with 4 decimal places.
@@
Finally, on the last attempt before our budget is exhausted, we preface the message with “This is the last try.” These prompts were not tuned for the task — we did not adjust them beyond adding “do not add anything else” to ensure that the language model output is easily parsable. Our reported results are a single seed with a temperature of 00 for the LLM.
B.3 Other Experiment Details
We use the default arguments for HyperparameterOptimizationFacade and BlackBoxFacade in the SMAC library (Lindauer et al., 2022), which correspond to using Random Forest and Gaussian Process surrogate models with BO.
B.4 Results
In Figure 8, we show the validation accuracy achieved on each task after 10 updates. In Table 5, we show results over trajectories of 60 iterations and 100 iterations. GPT-4 Turbo consistently beats random search and is comparable to Bayesian Optimization approaches.
Refer to captionRefer to caption
Figure 7: Minimum validation error achieved after $10$ function evaluations from different hyperparameter optimizers on datasets from HPOBench (Eggensperger et al., 2021). The benchmark defines a configuration search space on neural networks for learning rates, $\ell_{2}$ regularization, width, depth, and batch size. A large language model (GPT-4 Turbo) performs well compared to random search and Bayesian optimization ($BO$) with Gaussian process ($GP$) and random forest ($RF$) surrogate models. Aggregate results across four models are in Table 1.
Refer to captionRefer to caption
Figure 8: Minimum validation error achieved after $10$ function evaluations from different hyperparameter optimizers on datasets from HPOBench (Eggensperger et al., 2021). We show performance on each task tuning Random Forests and SVMs. Aggregate results across four models are in Table 1.
Table 5: Evaluating minimum validation error achieved for trajectories of length 60 and 100. We consider the SVMs, random forests, and logistic regression trained on eight datasets for a total of 24 tasks. Neural networks were too costly to train for trajectories of this length, so we used the other three models. GPT-4-Turbo consistently outperforms random search and is comparable to Bayesian optimization up to trajectories of length 100. The mean rank between all $4$ methods (each row + random) is $2.5$.
60 iterations
Model Beats Random ($\uparrow$) Median ($\uparrow$) Mean ($\uparrow$) Rank ($\downarrow$)
GPT-4 Turbo 87.50% 16.31% 22.76 % 1.94
$BO_{RF}$ 87.50% 19.61% 22.57 % 2.21
$BO_{GP}$ 91.67% 14.36% 24.78 % 2.19
100 iterations
Model Beats Random ($\uparrow$) Median ($\uparrow$) Mean ($\uparrow$) Rank ($\downarrow$)
GPT-4 Turbo 87.50% 11.71% 21.88 % 2.08
$BO_{RF}$ 91.67% 14.15% 19.69 % 2.33
$BO_{GP}$ 91.67% 12.75% 23.89 % 1.88
B.5 Additional Results
We consider the effect of appending a system prompt “You are a machine learning expert” in Table 6. We use this prompt throughout the rest of the experiments, following the idea of conditioning on good performance (Zhou et al., 2022).
Table 6: Evaluating effects of system prompt. We report the mean and median change as in the main paper.
Model Beats Random ($\uparrow$) Median ($\uparrow$) Mean ($\uparrow$)
Expert 90.62% 7.75% 23.23 %
Non-expert 81.25% 13.79% 22.36 %
Appendix C CIFAR-10 Experiment Details
We ran each experiment with three random seeds at temperature 0 for the LLM. We only compute the test accuracy on the 10000 points once, at the end of 20 training epochs. We use the compressed prompting approach with the following prompt:
You are helping tune hyperparameters for a neural network. This is our hyperparameter search space:
"optimizer": must be ["adam", "sgd"]
"learning_rate": between 1e-4 and 1e-1
"train_batch_size": 32, 64, 128, 256, 512
"weight_decay": between 1e-5 and 1e-1
"label_smoothing": between 0 and 0.2
You will get the validation error rate and loss before you need to specify the next configuration. The goal is to find the configuration that minimizes the error rate with the given budget, so you should explore different parts of the search space if the loss is not changing. Provide a config in JSON format. Do not put new lines or any extra characters in the response, only provide the config. Example config:
"optimizer": a
"learning_rate": b
"training batch size": c
"weight_decay": d
"label_smoothing": e
Random search samples from the same search space, with learning rate and weight decay sampled log-uniformly. We evaluate 100 samples and use bootstrapping to estimate the mean and standard error for plots. The unconstrained LLM prompt updates the first part of the prompt to:
"optimizer": must be ["adam", "sgd"]
"learning_rate": positive float
"train_batch_size": positive integer
"weight_decay": nonnegative float
"label_smoothing": nonnegative float
Appendix D Code Generation
@@
Appendix E 2-dimensional Landscape Experiment
To evaluate whether LLMs can reason about optimization choices, we study a set of 2-dimensional toy test functions commonly used in optimization: Rosenbrock (Rosenbrock, 1960), Branin (Dixon, 1978), Himmelblau (Himmelblau et al., 2018), and Ackley (Ackley, 1987). For each test function $f$, we consider optimizing $f(\boldsymbol{x})$ and also $f(\boldsymbol{x}-\boldsymbol{c})$, where $\boldsymbol{c}$ is a fixed constant, with $\boldsymbol{c}_{i}\sim\mathcal{U}(0,1)$, to mitigate the LLM’s potential to memorize the original $f$. Also, we benchmark performance on a well-conditioned and ill-conditioned quadratic. We provide the experimental details and the explored prompts in Appendix E.
We search for optimal solutions using LLMs as an optimizer with GPT-4. While these problems can be considered black-box optimization, we can also pose them as HPO to the LLM, using the same prompts in Figure 1 for consistency. We show the performance of a single fixed seed and temperature 00 in Figure 9. Performance across $3$ seeds outperforms random search in most tasks, as in Table 8. Moreover, inspecting the trajectories and reasoning chains shown in Appendix G, we qualitatively observe different behaviors that may characterize a strong search algorithm: initial exploration around search space boundary and center; a line-search type algorithm to probe function behavior; and reasoning that trades off between exploration and selecting a performant final value. However, performance can be inconsistent between random seeds, as shown in Table 7.
We use prompts with minimal details to gain intuition on this HPO approach and evaluate our LLM approach on problems with less complexity. Inspecting the trajectories, we see room for more specific prompts. For example, the LLMs re-explored previously evaluated points. We obtain the same or better performance in $8$ of the $10$ tasks by adding a sentence stating that the functions are deterministic, shown in Table 9. We speculate that HPO performance can generally be improved with more problem-setting details in the prompt.
Figure 9: The LLM optimizer optimized the function in most situations with few function evaluations, even when we apply random argument shifts to mitigate memorization. Darker colors represent regions of lower (better) loss, with the optimum denoted by a star.
E.1 Prompts
@@
Prompt 2 is similar to the second but frames the problem as hyperparameter optimization.
@@
We also study Prompt 3, adding a sentence: “The training process is deterministic and yields a nonnegative loss.”
Toy Function Prompt 0 Prompt 1 Prompt 2 Temp 0
@@
Table 7: Minimum loss achieved in $10$ iterations on $3$ different seeds at temperature $0.1$. We also run the experiments for Prompt 2 with temperature 00. Note that there is randomness in the results even at temperature 00.
Toy Function Prompt 0 Prompt 1 Prompt 2 Temp 00 Random
@@
Table 9: Average minimum loss achieved in 10 iterations on 3 different seeds at temperature 0. Note that there is randomness in the results even at temperature 0. Prompt 3 adds a sentence “The training process is deterministic and yields a nonnegative loss.” to Prompt 2. This achieves the same or better performance on $8$ of the $10$ tasks.
Appendix F LLMs as Tuning Assistants
@@
Figure 11: Response to question about low accuracies. The LLM-generated answer assumes the user is using a neural network and also suggests using dropout, which generally does not help for underfitting models.
Appendix G Trajectories
G.1 2D experiments
We provide code for loading the trajectories in load_toy_trajectories.
G.2 HPOBench
Here, we show some of the initial chain-of-thought reasoning for tuning neural networks on HPOBench generated by GPT-4.